Please help me with this investment problem!!
How much do I need to invest now at 8% compounded quarterly if I want the accumulated value of the investment to be $20,000 in 12 years?

Answer by checkley77(12844)

X(1+.08/4)^12*4=20,000
X*1.02^48=20,000
2.587X=20,000
X=20,000/2.587
X=$7,730.96 IS THE AMOUNT TO BE INVESTED TO ACHIEVE YOUR GOAL.